# Learn To be Efficient: Build Structured Sparsity in Large Language Models

### v3

@@ Abstract
Large Language Models (LLMs) have achieved remarkable success with their billion-level parameters, yet they incur high inference overheads. The emergence of activation sparsity in LLMs provides a natural approach to reduce this cost by involving only parts of the parameters for inference. However, existing methods only focus on utilizing this naturally formed activation sparsity in a post-training setting, overlooking the potential for further amplifying this inherent sparsity. In this paper, we hypothesize that LLMs can learn to be efficient by achieving more structured activation sparsity. To achieve this, we introduce a novel training algorithm, Learn-To-be-Efficient (LTE), designed to train efficiency-aware LLMs to learn to activate fewer neurons and achieve a better trade-off between sparsity and performance. Furthermore, unlike SOTA MoEfication methods, which mainly focus on ReLU-based models, LTE can also be applied to LLMs like LLaMA using non-ReLU activations. Extensive evaluation on language understanding, language generation, and instruction tuning tasks show that LTE consistently outperforms SOTA baselines. Along with our hardware-aware custom kernel implementation, LTE reduces LLaMA2-7B inference latency by 25% at 50% sparsity.
@@ 1 Introduction
Figure 1: Learn-To-be-Efficient (LTE) performs efficiency-aware training to construct more structured model contextual sparsity for fast inference. Activated neurons in FFN are in orange.
Large language models (LLMs), like GPT-3 [2], OPT [42], and LLaMA [35, 36], have demonstrated impressive natural language ability. However, the skyrocketing number of model parameters has not only inflated the deployment costs of LLMs, due to their significant computational and memory demands during inference [20], but also affected the user experience in many latency-sensitive applications, such as chatbots [1] and autonomous driving [16]. Recent advances have been leveraging sparsity to improve LLM inference efficiency, including model weight quantization [8] and pruning [21], token sparsity [44, 41], and activation sparsity [20, 43].
@@ 4.1 Experts Grouping
Following the previous work [43], we group 32 neurons as an expert in FFN layers, and use parameter clustering to group MLP neurons into different experts. Fundamentally, in FFN layers, each neuron is associated with a column of $W_{1}$ (which is a $d_{model}$ vector). Parameter clustering first treats the corresponding vector in $W_{1}$ as the feature for a neuron, then applies the balanced K-Means algorithm [22] to cluster neurons into $n$ clusters. Each cluster has $32$ neurons by default and will be treated as an individual MoE expert.111 For FFN layers using SwiGLU (e.g., LLaMA), there are two linear networks to calculate neuron values: $SwiGLU(x)=(Swish(xW)\otimes xV)W_{2}.$ We use the weight of the gate network, $W$, to cluster neurons. We also conduct an ablation study on other alternative grouping strategies in Section 5.4.
@@ 5.2 End-to-end Performance Comparison
50% FLOPs saving on NLG tasks. We next evaluate LTE with decoder-based models on three types of generation tasks. Similar to NLU tasks, Figure 6 shows that LTE outperforms all other baselines. NLG tasks are more challenging: all methods start to have a performance drop at a lower sparsity level
@@
compared to NLU tasks. Yet, LTE-trained models have better resilience to the sparsity increase. Especially at a high sparsity level, while other baselines fail to generate meaningful content, LTE still generates content with a marginal quality drop. Similarly, we also observe the performance drop caused by replacing soft activation with ReLU.
@@ 6 Conclusion
Broader Impact
This paper presents the Learn-To-be-Efficient (LTE) algorithm to accelerate the LLM inference. A more efficient LLM inference can significantly enhance the accessibility and sustainability of LLMs. By reducing computational demands, it addresses environmental concerns through lower energy consumption and helps democratize access to advanced AI technologies. We believe that our work can better help smaller organizations, educational institutions, and researchers, who previously faced barriers due to resource limitations, access LLMs more easily. To sum up, the LTE algorithm not only advances the field technically but also broadens the scope of AI’s benefits to a wider, more inclusive community.
Limitations
While LTE outperforms other baselines, LTE needs further fine-tuning to get contextual sparse models. (Yet, training is a one-time investment that yields long-term benefits for LLM serving.) A future study is to combine LTE with other parameter-efficient fine-tuning techniques to reduce LTE training overhead. Another limitation is our computational resource, which prevents us from training LTE on pretraining or RLHF. As discussed in Section 5.2, we believe that training with a wider variety of data will further improve LTE performance. We leave this for future exploration.
@@ Appendix B Experiment Settings
We presented all training hyperparameters in Table 234, and 5. For hyperparameters presented in {}, we select the best hyperparameter for each task. We follow the settings in RoBERTa paper [19] to fine-tune RoBERTa on GLUE datasets. We set the coefficient for the separability loss ($\lambda$ in Equation 6) to be 0.5 for all stage 1 training. We use different $\eta$ to control the sparsity of trained models (Figure 11). Hardware: All models are trained and evaluated on A100, A40, and 3090Ti, depending on memory usage and availability.

### v4

@@ Abstract
Large Language Models (LLMs) have achieved remarkable success with their billion-level parameters, yet they incur high inference overheads. The emergence of activation sparsity in LLMs provides a natural approach to reduce this cost by involving only parts of the parameters for inference. However, existing methods only focus on utilizing this naturally formed activation sparsity in a post-training setting, overlooking the potential for further amplifying this inherent sparsity. In this paper, we hypothesize that LLMs can learn to be efficient by achieving more structured activation sparsity. To achieve this, we introduce a novel training algorithm, Learn-To-be-Efficient (LTE), designed to train efficiency-aware LLMs to learn to activate fewer neurons and achieve a better trade-off between sparsity and performance. Furthermore, unlike SOTA MoEfication methods, which mainly focus on ReLU-based models, LTE can also be applied to LLMs like LLaMA using non-ReLU activations. Extensive evaluation on language understanding, language generation, and instruction tuning tasks show that LTE consistently outperforms SOTA baselines. Along with our hardware-aware custom kernel implementation, LTE reduces LLaMA2-7B inference latency by 25% at 50% sparsity. We make our code publicly available at GitHub [^1] https://github.com/haizhongzheng/LTE.
@@ 1 Introduction
Figure 1: Learn-To-be-Efficient (LTE) performs efficiency-aware training to construct structured model contextual sparsity for fast inference. Activated neurons are in orange.
The exponential growth in data volumes and model sizes has catalyzed significant breakthroughs in large-scale models, enabling a wide range of applications [2, 47, 53, 48, 30]. Among them, large language models (LLMs), like GPT-3 [2], OPT [47], and LLaMA [38, 39], have demonstrated impressive natural language ability. However, the skyrocketing number of model parameters [21, 36] and dataset size [51, 45, 52, 44] has introduced challenges in further scaling those models. The exponential growth in model size has not only inflated the deployment costs of LLMs, due to their significant computational and memory demands during inference, but also affected the user experience in many latency-sensitive applications, such as chatbots [1] and autonomous driving [17, 35]. Recent advances have been leveraging sparsity to improve LLM inference efficiency, including model weight quantization [9] and pruning [22], token sparsity [50, 46], and activation sparsity [21, 49].
@@ 4.1 Experts Grouping
Following the previous work [49], we group 32 neurons as an expert in FFN layers, and use parameter clustering to group MLP neurons into different experts. Fundamentally, in FFN layers, each neuron is associated with a column of $W_{1}$ (which is a $d_{model}$ vector). Parameter clustering first treats the corresponding vector in $W_{1}$ as the feature for a neuron, then applies the balanced K-Means algorithm [23] to cluster neurons into $n$ clusters. Each cluster has $32$ neurons by default and will be treated as an individual MoE expert.222 For FFN layers using SwiGLU (e.g., LLaMA), there are two linear networks to calculate neuron values: $SwiGLU(x)=(Swish(xW)\otimes xV)W_{2}.$ We use the weight of the gate network, $W$, to cluster neurons. We also conduct an ablation study on other alternative grouping strategies in Section 5.5.
@@ 5.2 End-to-end Performance Comparison
50% FLOPs saving on NLG tasks. We next evaluate LTE with decoder-based models on three types of generation tasks. Similar to NLU tasks, Figure 6 shows that LTE outperforms all other baselines. NLG tasks are more challenging: all methods start to have a performance drop at a lower sparsity level compared to NLU tasks. Yet, LTE-trained models have better resilience to the sparsity increase. Especially at a high sparsity level, while other baselines fail to generate meaningful content, LTE still generates content with a marginal quality drop. Similarly, we also observe the performance drop caused by replacing soft activation with ReLU.
@@ 5.3 Translate Sparsity to Wall-clock Time Speedup
5.4 Addtional Comparison
Model pruning. As discussed in Section 2, besides dynamic contextual sparsity provided by LTE, static sparsity provided by model pruning is also a common method to accelerate model inference. In this section, we present a performance comparison between Wanda [36] and LTE in Table 2. The sparsity reported here is the overall sparsity of the entire model, which is different from the FFN sparsity reported in Seciton 5.2. For LTE, we recalculate the overall sparsity based on the FFN sparsity. The evaluation results show that LTE achieves a lower perplexity than Wanda. The evaluation results show that, given a similar sparsity, LTE achieves better perplexity.
Table 2: Performance comparison with model pruning. We apply Wanda [36] on WikiText103 fine-tuned LLaMA-7B with the author-implemented code. Besides using C4 as the calibration data suggested by the Wanda paper, we also try to use WikiText for calibration to improve this baseline. The sparsity reported here is the overall sparsity of the entire model, which is different from the FFN sparsity. The evaluation results show that LTE achieves a lower perplexity than Wanda.
Method
Wanda
(2:4)
Wanda
(4:8)
Wanda
Wanda
(2:4)
Wanda
(4:8)
Wanda
LTE
($\eta=2$)
Overall Sparsity 50% 50% 50% 50% 50% 50% 52%
Calibration Data C4 C4 C4 Wiki Wiki Wiki -
PPL 11.45 8.39 7.04 10.82 8.17 6.87 5.95
@@ 5.5 Ablation Study
Figure 13: Performance comparison with new baselines: 1) $\sigma$-MoE [4] and 2) Deja VU [21] with further fine-tuning. We evaluate two baselines by fine-tuning GPT2 medium with WikiText103. LTE outperforms the other two baselines across different sparsity.
@@
Other baselines with further fine-tuning. A key distinction between LTE and other baselines is that LTE requires additional fine-tuning to make the model efficiency-aware. In this section, we assess whether further fine-tuning can enhance the performance of other baselines, aiming to understand the contribution of additional training to LTE’s effectiveness. Specifically, we fine-tune two baselines on GPT2-Medium using the WikiText103 dataset: 1) we first fine-tune GPT2-M with Wikitext and apply Deja Vu on the fine-tuned models. Then, we further fine-tune MoEfied models with WikiText103. 2) We implement $\sigma$-MoE [4] on GPT2-Medium model and fine-tune models with the WikiText103 dataset. For both baselines, we train models for the same training time as we train LTE models. The evaluation results shown in Figure 13 show that LTE outperforms the other two baselines across different sparsity levels.
@@ 6 Conclusion
Acknowledgement
My work was partially supported by DARPA Grant HR00112020008, National Science Foundation Grant No. 2039445, and Cisco Research Grant. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of our research sponsors.
@@ Appendix B Experiment Settings
We presented all training hyperparameters in Table 345, and 6. For hyperparameters presented in {}, we select the best hyperparameter for each task. We follow the settings in RoBERTa paper [20] to fine-tune RoBERTa on GLUE datasets. We set the coefficient for the separability loss ($\lambda$ in Equation 6) to be 0.5 for all stage 1 training. We use different $\eta$ to control the sparsity of trained models (Figure 11). Hardware: All models are trained and evaluated on A100, A40, and 3090Ti, depending on memory usage and availability.